OpenStudy (anonymous):

You need 1 1/4 cups of sugar to make 20 cookies To make 6 cookies, you will need cups of sugar. Enter your answer as a fraction or whole number.

OpenStudy (richyw):

\[\frac{5/4 \text{ cups}}{20\text{ cookies}}=\frac{1}{16}\text{cups/cookie}\]

OpenStudy (richyw):

does that make sense?

OpenStudy (anonymous):

I put 1/16 and it told me its wrong

OpenStudy (richyw):

well that is correct but it's not the answer you need. That tells you how many cups of sugar it takes PER COOKIE.

OpenStudy (richyw):

but how many cookies are you trying to make?

OpenStudy (anonymous):

20 right

OpenStudy (richyw):

nope. have you read the whole question?

OpenStudy (anonymous):

6 than bro ?

OpenStudy (richyw):

yup. so you know how much sugar it takes to make one cooke. It takes 1/16 cups. so how much sugar does it take to make 6 cookies?

OpenStudy (anonymous):

1/96 ?

OpenStudy (richyw):

not quite. remember \[6=\frac{6}{1}\]so\[\frac{1}{16}\times 6 = \frac{1}{16}\times\frac{6}{1}\]

OpenStudy (anonymous):

so whats the answer

OpenStudy (richyw):

I just gave it to you.

OpenStudy (anonymous):

lol rich don't hate me but Iis it 6/1 ?

OpenStudy (richyw):

nope. you need to multiply that by \(\frac{1}{16}\)

OpenStudy (richyw):

the answer is\[\frac{1}{16}\text{cups/cookie}\times 6\text{ cookies}\]

OpenStudy (richyw):

when you got 1/96 you did\[\frac{1}{16}\text{cups/cookie}\div 6\text{ cookies}\]you need to multiply instead of divide. you know how to multiply fractions right?

OpenStudy (anonymous):

nope.

OpenStudy (richyw):

you multiply the numerators together and the denominators together. \[\frac{a}{b}\times\frac{c}{d}=\frac{a\times c}{b\times d}\]

OpenStudy (anonymous):

can you please do it for me and I will check how you got it my calculator isn't working

OpenStudy (richyw):

you shouldn't need a calculator.

OpenStudy (richyw):

for example \[\frac{1}{3}\times 5 =\frac{1}{3}\times\frac{5}{1}=\frac{1\times 5}{3\times 1}=\frac{5}{3}\]

OpenStudy (anonymous):

can you label it for me like the example and I can do it

OpenStudy (richyw):

for the example you need to do, as I have said before\[\frac{1}{16}\times 6\]

OpenStudy (anonymous):

16/6?

OpenStudy (richyw):

nope. did you look at the example I gave you?

OpenStudy (richyw):

maybe you should go ask your parents how to multiply fractions.

OpenStudy (anonymous):

6/16

OpenStudy (anonymous):

my parents don't know lmao

OpenStudy (anonymous):

is it 6/16 ?

OpenStudy (anonymous):

im doing what u told me dude

OpenStudy (richyw):

yeah and you got the answer

OpenStudy (anonymous):

its 6/16 ?

OpenStudy (richyw):

that's how many cups of sugar it takes to make 6 cookies

OpenStudy (anonymous):

dang im sorry im so dumb man thanks

OpenStudy (richyw):

you aren't dumb. you just need to actually do math for yourself instead of posting it on open study and looking for easy answers. You have asked 200+ questions here and answered none.

OpenStudy (anonymous):

cause honestly I hate math and Im just asking questions because I cant think right with it

OpenStudy (anonymous):

alright thanks bye